OpenStudy (anonymous):

Find the exact value of k that makes the antidifferntation formula true (7x+18)^4 = k(7x+18)^5 + C

OpenStudy (anonymous):

Have you learned the substitution rule yet? You can use u=7x+18.

OpenStudy (amistre64):

take the derivative of the right side and compare it to the left side ....

OpenStudy (anonymous):

@amistre64 the derivative of the right side is 35(7x+18)^4 so what do i do know?

OpenStudy (amistre64):

you forgot to keep the k

OpenStudy (anonymous):

@BluFoot no im not sure what you are talking about

OpenStudy (anonymous):

35k(7x+18)^4?

OpenStudy (anonymous):

OK that's fine, use amistre's tips then :)

OpenStudy (amistre64):

35k(7x+18)^4 now, this needs to be exactly like the left side; so what value of k do we need to use?

OpenStudy (anonymous):

1/35?

OpenStudy (amistre64):

(7x+18)^4 = 35k (7x+18)^4 ; when k=1/35 great job ;)

OpenStudy (anonymous):

thx!